a horizontal force of 0.80N is required to pull a 5kg block across a table top at a constant speed. with the block initially at rest, a 20g bullet fired horizontally into the block causes the block to slide 1.5m before coming to rest again. determine the speed v of the bullet where the bullet is assumed to be embedded in the block


1
Expert's answer
2021-06-18T14:20:45-0400
"F=\mu Mg\\0.8=(5)(9.8)\mu\\\mu=0.016"

"0.5mv^2=\mu (m+M)gd\\
0.5(0.02)v^2=0.016(5.02)(9.8)(1.5)\\
v=11\frac{m}{s}"